OpenStudy (waleed_imtiaz):

how to find the volume of a glass with the help of integration and with the use of ruler? please help me

OpenStudy (tkhunny):

A drawing, perhaps?

OpenStudy (waleed_imtiaz):

suppose I have a base curve and I rotate it along y-axis and it makes a glass.... then the vloume of that glass would be whaT???

OpenStudy (tkhunny):

Insufficient defintion. Still needs a better description. For starters, if you've a curve, and you spin it around an axis, you will get an area, not a volume.

OpenStudy (waleed_imtiaz):

okey then as u are saying,,,tell me how to calculate the area with integration ...what will be the equation of that curve and the limits?

OpenStudy (tkhunny):

I recommend you read up on Pappus' Theorem. Find an effective centroid and spin away. It woudl help a lot if we had a clue what you were trying to do.

OpenStudy (slaaibak):

tkhunny, if you spin a curve around an axis, it forms something which you can calculate the volume of. you're thinking of a surface integral. but, the surface is the outer part of a figure that has volume.

OpenStudy (slaaibak):

http://en.wikipedia.org/wiki/Shell_integration

OpenStudy (tkhunny):

@slaaibak That is incorrect. Spinning a curve produces an area. Spinning an area under a curve, not that produces a volume.

OpenStudy (slaaibak):

tk - I agree, but that's what he is trying to say.

OpenStudy (tkhunny):

Fair enough. Do you think I've made enough fuss about it that the OP will now say what is meant? "-)

OpenStudy (slaaibak):

haha, we can only hope.

OpenStudy (waleed_imtiaz):

Well, u guyz are saying right but could u guyz plz give me an example solved to find the volume ....

OpenStudy (waleed_imtiaz):

I have made an attempt to find the volume ... and i use the equation of the base curve and the limits as radius of the base of the glass along x-axis.

OpenStudy (tkhunny):

Typically, using "Shells", we have: Given f(x) defned on a region [a,b] with 0 <= a < b And, for the sake of simplicity, f(x) >= 0 on [a,b] \(\large 2\pi\int\limits_{a}^{b} x\cdot f(x)\;dx\) If we use "Shells", we don't have to worry about the concavity of the shape.

OpenStudy (tkhunny):

Useful approximations can be made by measuring 'x' and relating it to f(x) for various widths and locations that capture the interesting parts of the curve. Notice, if f(x) = 1, a constant of height 1 on [0,2], we have: \(2\pi\int\limits_{0}^{2}x\cdot 1\;dx = 2\pi\cdot \dfrac{2^{2}}{2} = 4\pi\), which is consistent with the volume of a right circular cone with radius 2 and height = 1.

OpenStudy (waleed_imtiaz):

 can I do like that ? and then from these two points I can caluculate the equation f(x)... and then integrate it with using limits 0-5 ...guide me If I am wrong?